# CASteer: Cross-Attention Steering for Controllable Concept Erasure

### v3

@@ H.1.1 Method
H.1.2 Qualitative results
Here, we provide qualitative results results on flipping different concepts. Fig.30,31,32 visualize results of switching concepts of “Snoopy” to “Winnie-the-Pooh”, “cat” to “giraffe” and “caterpillar” to “butterfly”. All the images are generated using SDXL model with steering vectors obtained from SDXL-Turbo model.
Figure 30: Switching between concepts of “Snoopy" and “Winnie-the-Pooh" using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
Figure 31: Switching between concepts of “cat" and “giraffe" using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
Figure 32: Switching between concepts of “caterpillar" and “butterfly" using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
H.2 Concept addition
@@ H.2.2 Qualitative results
Fig.30,31,32 visualize results of adding concepts of “apple", “hat", “clothes" and “happiness" using CASteer. Note that for diferent concepts, different values of $\alpha$ are optimal.
Figure 33: Adding concepts of “apple", “hat", “clothes" and “happiness" using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
H.3 Style Transfer
@@ H.3.1 Method
CASteer performs Style Transfer on real images as follows: we apply the reverse diffusion process following DDIM DBLP:conf/iclr/SongME21 for $t$ number of steps, where $1\leqslant t\leqslant T$. Then we denoise image back using CASteer (addition algorithm, see eq.12). $t$ controls the trade-off between loss of image details and intensity of style applied. On fig. 34, 35, 36, 37 we show results on style transfer to four different styles with varying $t$. Intensity=0.3=0.3 here means that $t=0.3T$. With such a process, we often get satisfying results with no major loss of details. However, when $t$ is high, the loss of image content occurs (see bottom lines of figures). This is due to the fact that the inversion is not sufficiently accurate.
@@
Figure 34: Examples of Style Transfer of real images into “anime" style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
@@
Figure 35: Examples of Style Transfer of real images into “origami" style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
@@
Figure 36: Examples of Style Transfer of real images into “Gothic Art" style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
@@
Figure 37: Examples of Style Transfer of real images into “Retro Art" style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
Appendix I Interpreting steering vectors
I.1 Steering vectors visualization
@@
In this section, we propose a way to interpret the meaning of steering vectors generated by CASteer. Suppose we have steering vectors generated for a concept $\mathbf{X}$ $\{ca^{\mathbf{X}}_{it}\},1\leqslant i\leqslant n,1\leqslant t\leqslant T$, where $n$ is the number of model layers and $T$ is a number of denoising steps performed for generating steering vectors. To interpret these vectors, we prompt the diffusion model with a placeholder prompt “X" and at each denoising step, we substitute outputs of the model’s CA layers with corresponding steering vectors. This makes the diffusion model be only conditioned on the information from the steering vectors, completely suppressing other information from the text prompt.
@@
Fig. 38 shows interpretations for the steering vectors of concepts “hat", “polka dot dress", “Snoopy", “angry", “happy", from top to bottom. Note that vectors for concepts “hat", “polka dot dress" and “Snoopy" were generated using prompt pair templates for concept deletion, i.e. pairs of the form (“fish with Snoopy", “fish"), (“a girl with a hat", “a girl") (see sec.  C.1), and this is reflected in generated images, as they show these concepts not alone, but in a form of a girl in a hat, a girl in a polka dot dress or a boy with a Snoopy. As for the last two concepts (“angry", “happy"), their steering vectors were generated using prompt pair templates for human-related concepts (see sec.  C.2), and they illustrate these concepts as is.
@@
We note that images of each concept exhibit common features, e.g. all the images of hats and polka dots feature only female persons, and images corresponding to “angry" and “happy" concepts have certain styles. We believe this reflects how diffusion models perceive different concepts, and that this interpretation technique can be used for unveiling the hidden representations of concepts inside the diffusion model, but we leave it to future work.
@@
Figure 38: Visualization of generations of the model conditioned only on steering vectors. Images in rows from top to down were generated using steering vectors for the concepts “hat", “polka dot dress", “Snoopy", “angry", “happy"
I.2 UMap on steering vectors
@@
We generate steering vectors for all vocabulary tokens of SDXL text encoders and apply UMap DBLP:journals/corr/abs-1802-03426 on these steering vectors. Fig. 39414042 show that structure emerges in the space of these steering vectors, similar to that of Word2Vec DBLP:journals/corr/abs-1301-3781, supporting the hypothesis that steering vectors carry the meaning of the desired concept. This is observed for all the layers in the model.
@@
Figure 39: Visualization of UMap applied to the steering vectors from the layer 17 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens
@@
Figure 40: Visualization of UMap applied to the steering vectors from the layer 17 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens
@@
Figure 41: Visualization of UMap applied to the steering vectors from the layer 35 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens
@@
Figure 42: Visualization of UMap applied to the steering vectors from the layer 35 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens
Appendix J SPM vs CASteer on adversarial prompts
@@
In this section, we present more qualitative examples of CASteer outperforming SPM DBLP:conf/cvpr/Lyu0HCJ00HD24 on adversarial prompts, i.e. prompts containing implicitly defined concepts. SD-1.4 is used as a backbone for both methods. We use prompts “A mouse from Disneyland" and “A girl with a mouse from Disneyland" to test erasing of concept of “Mickey" and prompts “A yellow Pokemon" and “A girl with a yellow Pokemon" to test erasing of concept of “Pikachu".
@@
Figure 43: Examples of 8 generated images from CASteer and SPM when prompted “A mouse from Disneyland". Top: generation of CASteer, Bottom: generations of SPM. We use the same diffusion hyperparameters and seeds when generating corresponding images from CASteer and SPM
Figure 44: Examples of 8 generated images from CASteer and SPM when prompted “A girl with a mouse from Disneyland". Top: generation of CASteer, Bottom: generations of SPM. We use the same diffusion hyperparameters and seeds when generating corresponding images from CASteer and SPM
@@
Figure 45: Examples of 8 generated images from CASteer and SPM when prompted “A yellow Pokemon". Top: generation of CASteer, Bottom: generations of SPM. We use the same diffusion hyperparameters and seeds when generating corresponding images from CASteer and SPM

### v4

@@ I.1.1 Method
I.1.2 Qualitative results
Here, we provide qualitative results results on flipping different concepts. Fig. 333435 visualize results of switching concepts of “Snoopy” to “Winnie-the-Pooh”, “cat” to “giraffe” and “caterpillar” to “butterfly”. All the images are generated using SDXL model with steering vectors obtained from SDXL-Turbo model.
Figure 33: Switching between concepts of “Snoopy” and “Winnie-the-Pooh” using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
Figure 34: Switching between concepts of “cat” and “giraffe” using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
Figure 35: Switching between concepts of “caterpillar” and “butterfly" using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
I.2 Concept addition
@@ I.2.2 Qualitative results
Fig. 333435 visualize results of adding concepts of “apple", “hat", “clothes” and “happiness” using CASteer. Note that for different concepts, different values of $\alpha$ are optimal.
Figure 36: Adding concepts of “apple”, “hat”, “clothes” and “happiness” using CASteer. Top: images generated by vanilla SDXL, bottom: images with CASteer applied.
I.3 Style Transfer
@@ I.3.1 Method
CASteer performs Style Transfer on real images as follows: we apply the reverse diffusion process following DDIM Song et al. (2021) for $t$ number of steps, where $1\leqslant t\leqslant T$. Then we denoise image back using CASteer (addition algorithm, see Eq. 12). $t$ controls the trade-off between loss of image details and intensity of style applied. In Fig. 37,  38,  39, 40 we show results on style transfer to four different styles with varying $t$. Intensity=0.3=0.3 here means that $t=0.3T$. With such a process, we often get satisfactory results with no major loss of details. However, when $t$ is high, the loss of image content occurs (see the bottom lines of figures). This is because the inversion is not sufficiently accurate.
@@
Figure 37: Examples of Style Transfer of real images into “anime” style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
@@
Figure 38: Examples of Style Transfer of real images into “origami” style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
@@
Figure 39: Examples of Style Transfer of real images into “Gothic Art” style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
@@
Figure 40: Examples of Style Transfer of real images into “Retro Art” style. From top to bottom: original image, style transfer applied with intensities from 0.1 to 0.7 with a step of 0.1
Appendix J Interpreting steering vectors
J.1 Steering vectors visualization
@@
In this section, we propose a way to interpret the meaning of steering vectors generated by CASteer. Suppose we have steering vectors generated for a concept $\mathbf{X}$ $\{ca^{\mathbf{X}}_{it}\},1\leqslant i\leqslant n,1\leqslant t\leqslant T$, where $n$ is the number of model layers and $T$ is a number of denoising steps performed for generating steering vectors. To interpret these vectors, we prompt the diffusion model with a placeholder prompt “X” and at each denoising step, we substitute outputs of the model’s CA layers with corresponding steering vectors. This makes the diffusion model be only conditioned on the information from the steering vectors, completely suppressing other information from the text prompt.
@@
Fig. 41 shows interpretations for the steering vectors of concepts “hat”, “polka dot dress”, “Snoopy”, “angry”, “happy”, from top to bottom. Note that vectors for concepts “hat”, “polka dot dress” and “Snoopy” were generated using prompt pair templates for concept deletion, i.e. pairs of the form (“fish with Snoopy”, “fish”), (“a girl with a hat”, “a girl”) (see Sec.  C.1), and this is reflected in generated images, as they show these concepts not alone, but in a form of a girl in a hat, a girl in a polka dot dress or a boy with a Snoopy. As for the last two concepts (“angry”, “happy”), their steering vectors were generated using prompt pair templates for human-related concepts (see Sec.  C.2), and they illustrate these concepts as is.
@@
We note that images of each concept exhibit common features, e.g. all the images of hats and polka dots feature only female persons, and images corresponding to “angry” and “happy” concepts have certain styles. We believe that this reflects how diffusion models perceive different concepts and that this interpretation technique can be used for unveiling the hidden representations of concepts inside the diffusion model, but we leave it to future work.
@@
Figure 41: Visualization of generations of the model conditioned only on steering vectors. Images in rows from top to down were generated using steering vectors for the concepts “hat”, “polka dot dress”, “Snoopy”, “angry”, “happy”.
J.2 UMap on steering vectors
@@
We generate steering vectors for all vocabulary tokens of SDXL text encoders and apply UMap McInnes and Healy (2018) on these steering vectors. Fig. 42444345 show that structure emerges in the space of these steering vectors, similar to that of Word2Vec Mikolov et al. (2013), supporting the hypothesis that steering vectors carry the meaning of the desired concept. This is observed for all the layers in the model.
@@
Figure 42: Visualization of UMap applied to the steering vectors from the layer 17 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens.
@@
Figure 43: Visualization of UMap applied to the steering vectors from the layer 17 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens.
@@
Figure 44: Visualization of UMap applied to the steering vectors from the layer 35 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens.
@@
Figure 45: Visualization of UMap applied to the steering vectors from the layer 35 of SDXL-Turbo formed by 3000 SDXL-Turbo vocabulary tokens.
Appendix K SPM and DoCo vs CASteer on adversarial prompts
@@
In this section, we present more qualitative examples of CASteer outperforming SPM Lyu et al. (2024) and DoCo Wu et al. (2025) on adversarial prompts, i.e. prompts containing implicitly defined concepts. SD-1.4 is used as a backbone for both methods. We use prompts “A mouse from Disneyland” and “A girl with a mouse from Disneyland” to test erasing of concept of “Mickey”, prompts “A yellow Pokemon” and “A girl with a yellow Pokemon” to test erasing of concept of “Pikachu”, and prompt “A picture of Iron Man actor” to test erasing of concept of “Robert Downey Jr”. We use the code provided by SPM and DoCo to produce checkpoints and images for testing removal of the concepts of “Mickey” and “Pikachu”, and use checkpoint provided by DoCO to generate images for testing removal of a concept of “Robert Downey Jr”.
We see that SPM and DoCo fail to substantially remove targeted concepts in many cases, while CASteer shows consistent performance.
@@
Figure 46: Examples of 8 generated images from CASteer and SPM when prompted “A mouse from Disneyland”. Top: generation of CASteer, Middle: generations of SPM, Bottom: generations of DoCo.
@@
We use the same diffusion hyperparameters and seeds when generating corresponding images from CASteer, SPM and DoCo.